OpenStudy (anonymous):

Please help, I don't understand any of this... An investment that yielded 5% simple interest came due at the end of the year. Dominic received a check for the total amount (interest plus principle) of 136.5 What was his original investment?

OpenStudy (anonymous):

@welshfella can you help me?

OpenStudy (anonymous):

or know anyone who can?

OpenStudy (welshfella):

Let the principal be x Amount at the end of 1 year = x+ 5% of x 5% as a decimal is 0.05 so plugging in the given values 136.5 = x + 0.05x 136.5 = 1.05x can you find the value of x?

OpenStudy (welshfella):

you need to solve that equation for x.

OpenStudy (anonymous):

Thank you

OpenStudy (welshfella):